OpenStudy (anonymous):

Kelly loves to eat chocolate pizza. Each square inch of chocolate is 30 calories. She orders a slice of chocolate pizza from a pizza with a diameter of 12 inches that is cut with a central angle of 40 degrees. How many calories does she eat?

OpenStudy (anonymous):

about 3,393 calories about 113 calories about 13 calories about 377 calories might help you out alittle

OpenStudy (anonymous):

Hey trix?

OpenStudy (anonymous):

about 377 calories

OpenStudy (anonymous):

\[ \frac{22\times 4 \times 30}{7} \]

Directrix (directrix):

Kelly is eating a slice of pizza that is 40/360 or 1/9 of the pizza. 1/9 = K/ [2 ir (6^2)] where K is the area of Kelly's piece of pizz. 1/9 = K/ (72 pi) 9K = 72 pi K = 8 pi square inches Multiply that by 30 (calories per square inch)

OpenStudy (anonymous):

Hey Direct you good with slopes?

OpenStudy (chriss):

about 377 calories the area is pi6^2 of the entire pizza... a 40 degree cut would give about 1/9th of the pizza pi6^2/9 is 12.566 multiply that by 30 and you get 376.99

OpenStudy (anonymous):

thanks everyone

OpenStudy (chriss):

you're welcome. I had to pull out some old knowledge and think a bit for that one lol.